HUBBLE SPACE TELESCOPE DAILY REPORT #2381 PERIOD COVERED: 0000Z (UTC) 05/26/99 - 0000Z (UTC) 05/27/99 Daily Status Report as of 147/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 7621 (WF/PC-2 Cycle 7 Supplemental Darks PT1/3) The WF/PC-2 was used to make supplemental dark calibration images. This set of calibrations is intended to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed STIS/CCD/MA1 7437 (Understanding High-Redshift and Starburst Galaxies: A UV Spectroscopic Survey of O-Stars in the SMC) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make ultraviolet spectrographic observations of galaxy AV15. The data from this survey, of importance in many areas of astrophysics, will immediately be made public as no proprietary period is requested. A wavelength calibration image was also taken. As described in HSTAR 7147 and yesterday's report, there was an FGS loss of lock near the end of this proposal. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.3 Completed WF/PC-2 8059 (POMS Test Proposal: WF-2 Targeted Parallel Archive Proposal) The WF/PC-2 was used to make observations in parallel with the prime science observations. The parallel opportunities available with WF/PC-2 in the neighborhood of bright galaxies are treated in a slightly different way from the normal pure parallels. Local group galaxies offer the opportunity for a closer look at young stellar populations. As described in HSTAR 7147 and yesterday's report, there was an FGS loss of lock near the end of this proposal. Otherwise, The observations were completed as planned, and no further anomalies were reported. 1.4 Completed STIS/MA1/MA2 7436 (A Complete Sample of Hot Post-AGB Stars in Globular Clusters) The Space Telescope Imaging Spectrograph (CCD/MA1/MA2) was used to make observations of the globular cluster NGC-6752. Recent observations with the Ultraviolet Imaging Telescope have provided a complete sample of the hot post asymptotic giant branch {post-AGB} population in fourteen globular clusters. A wavelength calibration image was also taken. The observations were completed as planned, and no anomalies were reported. 1.5 Completed WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed STIS/CCD 7943 (Transmission of Filtered Echelle Slits) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations of the well-established spectrophotometric standard star BD+75D325 in order to measure the transmission of the filtered Echelle slits. Measurements were made between 1200 A0 and 3000 A0. The observations were completed as planned, and no problems were reported. 1.7 Completed WF/PC-2 8290 (Weather, Moons, and Orbit of the Brown Dwarf Gl 229B) The WF/PC-2 was used to make observations of the brown dwarf star HD42581 (a.k.a., GI 229B). Gl 229B is the only unambiguous brown dwarf. It can be imaged with high astrometric and photometric accuracy only by HST. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Three Sets of STIS/CCD 8406 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation V) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations as part of a POMS proposal. This is a STIS non-scripted parallel Condition V proposal. The observations were completed as scheduled, and no anomalies were reported. 1.9 Completed WF/PC-2 7504 (Calibration of Nearby Type Ia Supernovae as Standard Candles: The Next Three-Year Step) The WF/PC-2 was used to make observations of parent galaxy NGC-4527 in order to determining its Cepheid distance. This is to continue the calibration of the Type 1a supernova as standard candles. The observations were completed as planned, and no porblems were reported. 1.10 Completed WF/PC-2 8053 (WF/PC-2 Cycle-7 Earth Flats -- Part II) The WF/PC-2 was used to take Earth flat calibration images. This proposal obtains sequences of Earth streak flats to construct high quality flat fields for the WF/PC-2 filter set. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed WF/PC-2 7408 (Parallax, Proper Motion, and Spectral Energy Distribution of an Isolated Old Neutron Star) The WF/PC-2 was used to study the parallax, proper motion, and spectral energy distribution of the old neutron star RXJ185635. The target is a bright soft X-ray source in the direction of the R CrA molecular cloud. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed WF/PC-2 7470 (A Snapshot Survey of Galactic Globular Clusters) The WF/PC-2 was used to make snapshot survey observations of the crowded center of the galactic globular cluster NGC-6256. The resulting color-magnitude diagram will include all stars down to about one magnitude below the main-sequence turnoff. The observations were completed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 9 Successful : 9 As described in HSTAR 7149, there was a loss of lock at 147/030936Z. A proposal, STIS 7760, which occurs during the period of the next report, was affected. 2.2 FHST Updates: Scheduled : 16 Successful : 16 2.3 Operations Notes: HST operations continue to be conducted from the SAMS back-up facility using the PRS ground system. GSFC Internal Simulation #4 (EVA-3) was run yesterday. The simulation continues today with the exercise of the HST Deploy. Per ROP SR-1, the SSR EDAC error counter was cleared four times. The STIS MCE-2 reset at 146/130610Z while the low voltage was on and while in an SAA interval. MAMA-2 recovery was performed by normal SMS commanding at 147/0129Z. The STIS flight software error counter was reset at 146/1310Z as directed by ROP NS-12. A TTR was written for non-recoverable losses of data from 147/0054Z-0102Z and from 147/0107Z-0132Z due to mis-configured White Sands data port assignments. The NSSC-1 status buffer was dumped and reset at 147/0214Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH